An electron in a hydrogen atom in its ground state absorbs 1.5 times as much energy as the minimum required for it to escape from the atom. Find the Wavelength of the emitted electron

Grade:11

1 Answers

Arun
4 years ago
 
excess energy = 6.8 eV = 6.8 * 1.6 * 10^-19 J = 1.09 * 10^-18 J
h/mc = wavelength
here it can also be written as wavelength = h/sqrt(2mKE)
where h = 6.626 * 10^-34
m = 9.1 * 10^-31 kg
KE = eV = 1.09 * 10^-18 J
so substitute these answers and you will get answer as 4.7 angstrom.